# Xing Huanga), Panpan Renb), Feng-Yu Wanga)

### v1

@@ Abstract
Let $\tilde{W}^{-\delta,k}$ be the local negative Sobolev space on $\mathbb{R}^{d}$ with indexes $\delta\in[0,\infty)$ and $k\in[1,\infty].$ We study McKean-Vlasov SDEs with interaction kernels in $\tilde{W}^{-\delta,k}.$ By developing a time-shift argument which allows the singular interactions vanishing at time 0, the global well-posedness is proved for regular enough initial distributions and any singular indexes $(\delta,k)\in[0,\infty)\times[1,\infty],$ and for any initial distributions provided $\delta+\frac{d}{k}<1$. Moreover, the relative entropy and the $\|\cdot\|_{\delta,k*}$-distance induced by $\tilde{W}^{-\delta,k}$ are estimated for the time-marginal distributions of solutions by using the Wasserstein distance of initial distributions, which describe the regularity of the solution in initial distribution. In particular, the main results apply to Nemytskii-type SDEs which point-wisely depend on the density function and its derivatives, as well as McKean-Vlasov SDEs with interactions more singular than Riesz kernels.
@@
Keywords: McKean-Vlasov SDE, local distributional interaction, $\|\cdot\|_{\delta,k*}$-distance,
entropy-cost inequality.
@@ 1 Introduction
for $B(z,1):=\{x\in\mathbb{R}^{d}:|x-z|\leq 1\}$ and $\|\cdot\|_{L^{k}}$ being the $L^{k}$-norm with respect to the Lebesgue measure. Let $\tilde{W}_{*}^{-\delta,k}$ be the dual space of $\tilde{W}^{-\delta,k}.$ We study (1.1) for $h_{t}\in\tilde{W}^{-\delta,k}$ for some $(\delta,k)\in[0,\infty)\times[1,\infty],$ where $b_{t}$ in (1.2) is well-defined as duality for
$\mu\in\mathscr{P}_{\delta,k*}:=\mathscr{P}\cap\tilde{W}_{*}^{-\delta,k}:=\left\{\mu\in\mathscr{P}:\ \|\mu\|_{\delta,k*}:=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}|\mu(f)|<\infty\right\}.$
@@ 3. $(3)$
When $\kappa=0$, the condition (2.1) coincides with [4, (1.33)] for $\alpha=2,q_{b}=\infty,p_{0}=\frac{p}{p-1},\rho_{0}=k,\beta_{0}=\varepsilon$ and $\beta_{b}=-\delta$. In this case, [4, Theorem 1.9] ensures the well-posedness and density estimates for (1.1) provided $\|h_{t}\|_{\mathbb{B}_{k,\infty}^{\delta}}<\infty$. If moreover $p=\infty$ and $\varepsilon=0$, (2.1) becomes $\delta<1-\frac{d}{k},$ so that [1, Theorem 1] with $\alpha=2$ implies the global well-posedness of (1.1) for $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty.$ These results do not cover Theorem 2.1 since $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty$ may fail for $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d}).$
@@
We present the following two examples to illustrate Theorem 2.1 and Theorem 2.3, where the kernel $h$ is more singular than the Riesz kernel as considered in previous papers, see [8, 11, 13, 14] and references therein, and covers the Dirac interaction for Nemytskii-type SDEs depending on the density and any order derivatives.
@@ 3 Some preparations
On the other hand, by [8, Lemma 5.3], there exists a constant $c_{0}\in(0,\infty)$ such that
@@ 2. $(2)$
Then the desired assertion follows from [8, Proposition 5.1]. ∎
@@
By [8, Proposition 5.4], (3.6) implies that for $p_{2}>1$
@@
Moreover, by the Duhamel formula, see [8, Proposition 5.5(2)], we have
@@
4 Proof of Theorem 2.1
@@ 3. $(3)$
According to the proof of [8, Theorem 2.1], the first assrtion follows from the second. So, it suffices to show that for any $n\in\mathbb{N}$ and initial value $X_{0}$ with $\gamma\in\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ (1.1) has a unique weak/strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution up to $\tau_{n}(\gamma)$ such that
@@
Then by (2.1) and (2.5), there exists $q^{\prime}>2$ such that $\|b^{\mu}\|_{\tilde{L}_{q^{\prime}}^{\infty}(\tau_{n})}$ in (3.6) is bounded above by a constant $D_{n,\gamma}\in(0,\infty)$ depending on $n$ and $\|\gamma\|_{\varepsilon,p*}$. So, (4.15) follows from [17, Theorem 1.3.1], where $c_{n,q}(\gamma)$ is uniform in $\gamma\in\mathscr{P}$ when $\varepsilon=0$ and $p=\infty$, since $\|\gamma\|_{0,\infty*}=1$. ∎
@@
for a nonnegative measurable function $f$. By [18, Lemma 2.1] and $P_{t}^{\gamma}f\in C(\mathbb{R}^{d})$ for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ due to (5.5), we find a constant $k_{1}\in(0,\infty)$ such that
@@
(b) By Duhamel’s formula [8, Proposition 5.5 (2)], we have
@@
The proof is similar to that of [8, Theorem 2.3(2)], where $\varepsilon=\delta=0.$ For completeness we figure out it in the present situation.
@@
So, by [8, Proposition 5.2(4)], we derive (5.13) for $0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma))$ and some $c_{0}(t)\in(0,\infty)$ increasingly in $t$. When $t-s>s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma)),$ by the semigroup property and Jensen’s inequality, we deduce
@@
Combining this with [8, Proposition 5.5(1)], we find $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that

### v2

@@ Abstract
Let $\tilde{W}^{-\delta,k}$ be the local negative Sobolev space on $\mathbb{R}^{d}$ with indexes $\delta\in[0,\infty)$ and $k\in[1,\infty].$ We study McKean-Vlasov SDEs with interaction kernels in $\tilde{W}^{-\delta,k}.$ By developing a time-shift argument which allows the singular interactions vanishing at time 0, the global well-posedness is proved for regular enough initial distributions and any singular indexes $(\delta,k)\in[0,\infty)\times[1,\infty],$ and for any initial distributions provided $\delta+\frac{d}{k}<1$. Moreover, the relative entropy and the $\|\cdot\|_{\delta,k*}$-distance induced by $\tilde{W}^{-\delta,k}$ are estimated for the time-marginal distributions of solutions by using the Wasserstein distance of initial distributions, which describe the regularity of the solution in initial distribution. In particular, the main results apply to Nemytskii-type SDEs depending on higher order derivatives of density functions, as well as McKean-Vlasov SDEs with interactions more singular than Riesz kernels.
@@
Keywords: McKean-Vlasov SDE, local distributional interaction, well-posedness,
$\|\cdot\|_{\delta,k*}$-distance, entropy-cost inequality.
@@ 1 Introduction
for $B(z,1):=\{x\in\mathbb{R}^{d}:|x-z|\leq 1\}$ and $\|\cdot\|_{L^{k}}$ being the $L^{k}$-norm with respect to the Lebesgue measure. Let $\tilde{W}_{*}^{-\delta,k}$ be the dual space of $\tilde{W}^{-\delta,k}.$ We study (1.1) for $h_{t}\in\tilde{W}^{-\delta,k}$ for some $(\delta,k)\in[0,\infty)\times[1,\infty],$ where $b_{t}$ in (1.2) is well-defined as duality for $\mu\in\mathscr{P}_{\delta,k*}:=\mathscr{P}\cap\tilde{W}_{*}^{-\delta,k}.$ We have
$\mathscr{P}_{\delta,k*}=\left\{\mu\in\mathscr{P}:\ \|\mu\|_{\delta,k*}:=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}|\mu(f)|<\infty\right\}.$
@@ 3. $(3)$
When $\kappa=0$, the condition (2.1) coincides with [6, (1.33)] for $\alpha=2,q_{b}=\infty,p_{0}=\frac{p}{p-1},\rho_{0}=k,\beta_{0}=\varepsilon$ and $\beta_{b}=-\delta$. In this case, [6, Theorem 1.9] ensures the well-posedness and density estimates for (1.1) provided $\|h_{t}\|_{\mathbb{B}_{k,\infty}^{\delta}}<\infty$. If moreover $p=\infty$ and $\varepsilon=0$, (2.1) becomes $\delta<1-\frac{d}{k},$ so that [2, Theorem 1] with $\alpha=2$ implies the global well-posedness of (1.1) for $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty.$ These results do not cover Theorem 2.1 since $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty$ may fail for $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d}).$
@@
We present the following two examples to illustrate Theorem 2.1 and Theorem 2.3, where the kernel $h$ is more singular than the Riesz kernel as considered in previous papers, see [5, 10, 13, 15, 16] and references therein. In particular, our results apply to the Dirac interaction for Nemytskii-type SDEs depending on the density and its higher order derivatives, which include the Burgers/Navier-Stokes/$p$-Laplacian equations as typical examples where the first and second order derivatives of density are involved in, see for instance [1, 17].
@@ 3 Some preparations
On the other hand, by [10, Lemma 5.3], there exists a constant $c_{0}\in(0,\infty)$ such that
@@ 2. $(2)$
Then the desired assertion follows from [10, Proposition 5.1]. ∎
@@
By [10, Proposition 5.4], (3.6) implies that for $p_{2}>1$
@@
Moreover, by the Duhamel formula, see [10, Proposition 5.5(2)], we have
@@
4 Proof of the existence and uniqueness
In this section we prove Theorem 2.1.
@@ 3. $(3)$
According to the proof of [10, Theorem 2.1], the first assrtion follows from the second. So, it suffices to show that for any $n\in\mathbb{N}$ and initial value $X_{0}$ with $\gamma\in\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ (1.1) has a unique weak/strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution up to $\tau_{n}(\gamma)$ such that
@@
Then by (2.1) and (2.5), there exists $q^{\prime}>2$ such that $\|b^{\mu}\|_{\tilde{L}_{q^{\prime}}^{\infty}(\tau_{n})}$ in (3.6) is bounded above by a constant $D_{n,\gamma}\in(0,\infty)$ depending on $n$ and $\|\gamma\|_{\varepsilon,p*}$. So, (4.15) follows from [20, Theorem 1.3.1], where $c_{n,q}(\gamma)$ is uniform in $\gamma\in\mathscr{P}$ when $\varepsilon=0$ and $p=\infty$, since $\|\gamma\|_{0,\infty*}=1$. ∎
@@ 5 Proof of regularity estimates
for a nonnegative measurable function $f$. By [21, Lemma 2.1] and $P_{t}^{\gamma}f\in C(\mathbb{R}^{d})$ for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ due to (5.5), we find a constant $k_{1}\in(0,\infty)$ such that
@@
(b) By Duhamel’s formula [10, Proposition 5.5 (2)], we have
@@
The proof is similar to that of [10, Theorem 2.3(2)], where $\varepsilon=\delta=0.$ For completeness we figure out it in the present situation.
@@
So, by [10, Proposition 5.2(4)], we derive (5.13) for $0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma))$ and some $c_{0}(t)\in(0,\infty)$ increasingly in $t$. When $t-s>s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma)),$ by the semigroup property and Jensen’s inequality, we deduce
@@
Combining this with [10, Proposition 5.5(1)], we find $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that
@@
Data Availability
Data sharing is not applicable to this paper as no datasets were generated or analyzed.
Declarations
The authors declare that they have no Conflict of interests.